# Improving Spoken Language Modeling with Phoneme Classification: A Simple Fine-tuning Approach

### v1

@@ Abstract
Recent progress in Spoken Language Modeling has demonstrated the feasibility of learning language directly from speech. Generating speech through a pipeline that operates at the text level typically loses nuances, intonations, and non-verbal vocalizations. Modeling directly from speech opens up the path to more natural and expressive systems. On the other hand, speech-only systems tend to trail behind text-based language models in terms of their semantic abilities. We show that fine-tuning speech representation models on phoneme classification leads to more context-invariant representations, which in turn improve downstream language modeling performance.
@@ 1 Introduction and related work
Recent advances in Self-supervised Speech Representation Learning (SSL) Mohamed et al. (2022); Chen et al. (2022); Hsu et al. (2021); Baevski et al. (2020) have enabled the development of label-free representations that are valuable for various downstream tasks wen Yang et al. (2021). These representations can be discretized and treated as pseudo-text, allowing for the training of language models directly from raw audio Lakhotia et al. (2021), which capture both prosody and linguistic content Kharitonov et al. (2022). Applications of these audio-based language models include dialogue modeling Nguyen et al. (2023b), emotion conversion Polyak et al. (2021), and direct speech-to-speech translation Lee et al. (2022). They can be trained not only on discretized SSL representations but also on continuous word-size tokens Algayres et al. (2023) or a combination of acoustic and semantic tokens Borsos et al. (2023). However, these models still lag behind their text-based counterparts in terms of capturing semantics when trained with similar data quantity Nguyen et al. (2020). Recent approaches tackled this issue by jointly training speech and text Language Models (LMs) Nguyen et al. (2024); Maiti et al. (2024); Chou et al. (2023) or by using existing LMs as a warm initialization Hassid et al. (2023).
Unfortunately, SSL units are not invariant to irrelevant acoustic variations, which hinders downstream language modeling. Recent works have addressed this issue for background noise Chen et al. (2022), speech rate change Gat et al. (2023), and speaker change Qian et al. (2022); Chang et al. (2023); Chang and Glass (2023). However, contextual variations due to coarticulation remain a challenge Hallap et al. (2023): SSL units align more closely with contextual phone states Young et al. (1994) rather than linguistic units Dunbar et al. (2022), which may affect the LM’s capacity to learn higher-order representations of language.
Here, we test a simple idea: using supervised fine-tuning from a phoneme classification task to help the model remove its contextual dependency. We first show that fine-tuned models learn representations that are much more context-invariant than the original SSL representations, even with as little as a few hours of labels. Next, we show that these representations can be used to train an LM that outperforms the standard approach. We then evaluate whether the fine-tuned representations have retained their expressive power by measuring the distortion when resynthesizing expressive speech.
@@ 2.1 Phoneme classification
We started from the pretrained HuBERT Hsu et al. (2021) Base model, with 95M parameters, and fine-tuned it on a phoneme classification task. Instead of using a CTC objective Graves et al. (2006), we resorted to frame-level phoneme classification. Our main focus is not the quality of the phonemic transcription of the audio. Our goal is that individual representations are well aligned with the phonemic content of the audio source. By operating at the frame level, we impose a strong condition on each frame and do not lose the temporal component. Such an approach comes at the cost of needing forced alignments from the training dataset, when using CTC only demands sentence-level transcripts. We added one fully connected layer on top of the HuBERT backbone that maps the 768-dimensional representation to our phoneme space of dimension 40. We fine-tuned this model on LibriSpeech train-clean-100 Panayotov et al. (2015). We also reported results for models fine-tuned on LibriLight Limited , , and Kahn et al. (2020). The forced alignments are those used in Nguyen et al. (2020), obtained with the Abkhazia library111https://github.com/bootphon/abkhazia. The fine-tuning hyperparameters are derived from those used in Hsu et al. (2021) for ASR. We trained for steps with a batch size of 32 on a single NVIDIA V100 GPU.
@@ 2.4 Speech resynthesis
For speech resynthesis, we trained a HiFi-GAN Kong et al. (2020); Polyak et al. (2021) on the Expresso dataset Nguyen et al. (2023a), conditioned on the HuBERT discrete speech units and one-hot speaker embeddings from one of Expresso’s voices. We trained for steps on two NVIDIA V100 GPUs and followed the other hyperparameters used in Expresso. In this setup the HiFi-GAN has a different training domain than the HuBERT, the k-means, and the LM, which were trained on the audiobooks of LibriSpeech.
@@ 2.5 Evaluation metrics
Following Hallap et al. (2023), we also evaluate our models on the Phoneme ABX task, where each token is a phoneme. We examine two conditions: “within context” (constant preceding and following phonemes) and “any context” (no constraints on context). This task assesses context-invariance in speech representations, revealing that current self-supervised systems struggle with context independence. Notably, in Hallap et al. (2023) the performance drop when removing the constant context condition is larger than the gaps observed in speaker independence or clean versus less-clean speech conditions. By fine-tuning at a frame level without taking into account the context, our approach is a way to directly tackle this issue.
@@ 3.2 Results above the phonemic level
We report in Table 3 the zero-shot sWUGGY (lexical level) and sBLIMP (syntactic level) scores for the base and fine-tuned models, as well as for an LSTM trained on the gold phonemes. Following the observation regarding the ABX error rates of the centroids, which remained within 1 percentage point of the standard continuous units, we train LSTMs by initializing their embedding table directly with the associated centroid representation of dimension 768. Apart from this change, the training process is the same between the two conditions. As can be seen, fine-tuning for phoneme classification improves spoken language modeling in terms of zero-shot comprehension evaluations. Overall, the gap between training from speech and training with golden phonemes is now halved. Fine-tuning for phoneme classification results in models that are on par in terms of lexical comprehension with much larger baselines, which were trained on orders of magnitude more of data.
@@ 4 Conclusion
We showed that fine-tuning SSL representations with a phoneme classification task is an effective and simple procedure to improve context independence. This leads to improvements in the performance of LMs trained on these units. And we also found that initializing the embeddings of the discrete tokens of the LMs with the centroids of the units further helps with LM scores. This shows that the units found are meaningfully placed relative to one another in this representation space. Our work also highlights the trade-off between language modeling (which works best with abstract units), and speech generation (which works best with specific units). Fine-tuning on phoneme classification can adjust this trade-off.
@@ 5 Limitations
Further work is needed to improve on the trade-off, perhaps by combining SSL, resynthesis, and fine-tuning objectives concurrently. More comprehensive studies could explore the role of the encoder in the spoken language modeling pipeline by examining the impact of fine-tuning methods on downstream language modeling, comparing various SSL and supervised speech models. Another important direction to consider is the application of this method in a multilingual setting. The benefits of fine-tuning are visible after training on as little as a few hours of aligned data, making it applicable to low resource languages.
@@ Appendix A Appendix
A.1 Fine Tuning Results

### v2

@@ Abstract
Recent progress in Spoken Language Modeling has shown that learning language directly from speech is feasible. Generating speech through a pipeline that operates at the text level typically loses nuances, intonations, and non-verbal vocalizations. Modeling directly from speech opens up the path to more natural and expressive systems. On the other hand, speech-only systems require up to three orders of magnitude more data to catch up to their text-based counterparts in terms of their semantic abilities. We show that fine-tuning speech representation models on phoneme classification leads to more context-invariant representations, and language models trained on these units achieve comparable lexical comprehension to ones trained on hundred times more data.
@@ 1 Introduction and related work
Recent advances in Self-supervised Speech Representation Learning (SSL) Mohamed et al. (2022); Chen et al. (2022); Hsu et al. (2021); Baevski et al. (2020) have enabled the development of label-free representations that are valuable for various downstream tasks wen Yang et al. (2021). These representations can be discretized and treated as pseudo-text, allowing for the training of language models directly from raw audio Lakhotia et al. (2021), which capture both prosody and linguistic content Kharitonov et al. (2022). Applications of these audio-based language models include dialogue modeling Nguyen et al. (2023b), emotion conversion Polyak et al. (2021), and direct speech-to-speech translation Lee et al. (2022). They can be trained not only on discretized SSL representations but also on continuous word-size tokens Algayres et al. (2023) or on a combination of acoustic and semantic tokens Borsos et al. (2023). However, these models still lag behind their text-based counterparts in terms of capturing semantics when trained with similar data quantity Nguyen et al. (2020), with scaling laws up to three orders of magnitude slower Cuervo and Marxer (2024). Recent approaches tackled this issue by jointly training speech and text Language Models (LMs) Nguyen et al. (2024); Maiti et al. (2024); Chou et al. (2023) or by using existing LMs as a warm initialization Hassid et al. (2023).
One hypothesis for the data inefficiency of spoken language models is that they must at the same time perform language modeling and process irrelevant acoustic variations. Recent works have addressed this issue for background noise Chen et al. (2022), speech rate change Gat et al. (2023), and speaker change Qian et al. (2022); Chang et al. (2023); Chang and Glass (2024). However, contextual variations due to coarticulation remain a challenge Hallap et al. (2023): SSL units align more closely with contextual phone states Young et al. (1994) than with linguistic units Dunbar et al. (2022), which may affect the LM’s capacity to learn higher-order representations of language.
Here, we test a simple idea: using supervised fine-tuning on a phoneme classification task to help the model remove its contextual dependency. We first show that fine-tuned models learn representations that are much more context-invariant than the original SSL representations, even with as little as a few hours of labels. Next, we show that these representations can be used to train a LM that outperforms the standard approach. We then evaluate whether the fine-tuned representations have retained their expressive power by measuring the distortion when resynthesizing expressive speech.
We release the code and models at https://github.com/bootphon/spokenlm-phoneme.
@@ 2.1 Phoneme classification
We started from the pretrained HuBERT Hsu et al. (2021) Base model, with 95M parameters, and fine-tuned it on a frame-wise phoneme classification task with a forced aligned gold transcription. We chose this objective to give the model full information about phoneme identity and boundaries, to enforce the learning of context-invariant representations. An alternative would have been to use a CTC objective Graves et al. (2006), which has the advantage of not requiring forced-alignment, but may result in alignment errors hindering context-invariance. As shown in Appendix A.1, CTC fine-tuning results in slightly lower performance than phone classification.
We added one fully connected layer on top of the HuBERT backbone that maps the 768-dimensional representation to our phoneme space of dimension 40. We fine-tuned this model on LibriSpeech train-clean-100 Panayotov et al. (2015). We also reported results for models fine-tuned on LibriLight Limited , , and Kahn et al. (2020). The forced alignments are those used in Nguyen et al. (2020), obtained with the Abkhazia library111https://github.com/bootphon/abkhazia. The fine-tuning hyperparameters are derived from those used in Hsu et al. (2021) for ASR.
Input frames are partially masked as in pretraining, but the prediction loss is computed over all output frames, not just the masked ones. We trained for steps with a batch size of 32 on a single NVIDIA V100 GPU.
@@ 2.4 Speech resynthesis
For speech resynthesis, we trained a HiFi-GAN Kong et al. (2020); Polyak et al. (2021) on the Expresso dataset Nguyen et al. (2023a), conditioned on the HuBERT discrete speech units and one-hot speaker embeddings from one of Expresso’s voices. We trained for steps on two NVIDIA V100 GPUs and followed the other hyperparameters used in Expresso. In this setup the HiFi-GAN has a different training domain than the HuBERT, the k-means, and the LM, which were trained on the audiobooks of LibriSpeech. Expresso is rich in expressive variations, paralinguistics and nonvocals, making it well-suited to evaluate whether the discrete units preserve expressivity along with phonemic content.
@@ 2.5 Evaluation metrics
Following Hallap et al. (2023), we also evaluate our models on the Phoneme ABX task, where each token is a phoneme. We examine two conditions: “within context” (constant preceding and following phonemes) and “any context” (no constraints on context). This task assesses context-invariance in speech representations, revealing that current self-supervised systems struggle with context independence. Notably, in Hallap et al. (2023) the performance drop when removing the constant context condition is larger than the gaps observed in speaker independence or clean versus less-clean speech conditions. By fine-tuning at a frame level without taking into account the context, our approach is a way to directly tackle this issue. For complementary analysis of the discrete units, see Appendix A.2.
@@ 3.2 Results above the phonemic level
We report in Table 3 the zero-shot sWUGGY (lexical level) and sBLIMP (syntactic level) scores for the base and fine-tuned models, as well as for an LSTM trained on the gold phonemes. Following the observation regarding the ABX error rates of the centroids, which remained within 1 percentage point of the standard continuous units, we train LSTMs by initializing their embedding table directly with the associated centroid representation of dimension 768. Apart from this change, the training process is the same between the two conditions. Fine-tuning for phoneme classification improves spoken language modeling in terms of zero-shot comprehension evaluations. Overall, the gap between training from speech and training with golden phonemes is now halved. Fine-tuning for phoneme classification results in models that are on par in terms of lexical comprehension with much larger baselines, which were trained on orders of magnitude more of data.
@@ 4 Conclusion
We showed that fine-tuning SSL representations with a phoneme classification task is an effective and simple procedure to improve context independence. LMs trained on these units achieve comparable lexical comprehension to models trained on hundred times more data. And we also found that initializing the embeddings of the discrete tokens of the LMs with the centroids of the units further helps with LM scores. This shows that the units found are meaningfully placed relative to one another in this representation space. Our work also highlights the trade-off between language modeling (which works best with abstract units), and speech generation (which works best with specific units). Fine-tuning on phoneme classification can adjust this trade-off.
@@ 5 Limitations
Further work is needed to improve on the trade-off, perhaps by combining SSL, resynthesis, and fine-tuning objectives concurrently. More comprehensive studies could explore the role of the encoder in the spoken language modeling pipeline by examining the impact of fine-tuning methods on downstream language modeling, comparing self-supervised and supervised speech models with different kinds of supervision. Another important direction to consider is the application of this method in a multilingual setting. The benefits of fine-tuning are visible after training on as little as a few hours of aligned data, making it applicable to low resource languages.
@@ Appendix A Appendix
A.1 Fine-tuning results
As preliminary investigation, we fine-tuned HuBERT on phoneme recognition with a CTC loss instead of frame-level classification. As shown in Figure 3 this results in slightly weaker performance in terms of ABX.